OpenStudy (anonymous):

The expression 1000(1.1)^t represents the value of a $1000 investment that earns 10% interest per year, compunded annually for t years. What is the value of a $1000 investment at the end of each period? 1) 2 years 2) 4 years

OpenStudy (nurali):

For 2 Year: A=1000(1.1)^2 A=$1210 For 4 Year: A=1000(1.1)^4 A=$1464.10

OpenStudy (nurali):

Welcome to Openstudy.

OpenStudy (anonymous):

Thank you! :)

OpenStudy (nurali):

Anytime